### v4

# Heterogeneous Multi-Agent Proximal Policy Optimization for Power Distribution System Restoration

Abstract
Restoring power distribution systems (PDS) after large-scale outages requires sequential switching operations that reconfigure feeder topology and coordinate distributed energy resources (DERs) under nonlinear constraints such as power balance, voltage limits, and thermal ratings. These challenges make conventional optimization and value-based RL approaches computationally inefficient and difficult to scale. This paper applies a Heterogeneous-Agent Reinforcement Learning (HARL) framework, instantiated through Heterogeneous-Agent Proximal Policy Optimization (HAPPO), to enable coordinated restoration across interconnected microgrids. Each agent controls a distinct microgrid with different loads, DER capacities, and switch counts, introducing practical structural heterogeneity. Decentralized actor policies are trained with a centralized critic to compute advantage values for stable on-policy updates. A physics-informed OpenDSS environment provides full power-flow feedback and enforces operational limits via differentiable penalty signals rather than invalid action masking. The total DER generation is capped at 2400 kW, and each microgrid must satisfy local supply–demand feasibility. Experiments on the IEEE 123-bus and IEEE 8500-node systems show that HAPPO achieves faster convergence, higher restored power, and smoother multi-seed training than DQN, PPO, MAES, MAGDPG, MADQN, Mean-Field RL, and QMIX. Results demonstrate that incorporating microgrid-level heterogeneity within the HARL framework yields a scalable, stable, and constraint-aware solution for complex PDS restoration.
I Introduction
Modern power distribution systems face increasing exposure to environmental hazards, component aging, and cybersecurity threats, resulting in frequent service interruptions [1]. Distributed energy resources (DERs) such as photovoltaics, microturbines, and storage can accelerate local recovery [2], yet their limited capacity and the voltage, thermal, and topological constraints of distribution feeders complicate safe and efficient restoration [3]. Recent advancements in machine learning for critical infrastructure highlight the growing role of data-driven recovery, including a transformer–reinforcement learning framework for large-scale repair sequencing that enables faster and more resilient system restoration [4]. Machine learning has also advanced modeling in other safety-critical physical domains, including high-fidelity ground-motion prediction using ensemble techniques [5].
Conventional restoration is typically formulated as a mixed-integer nonlinear optimization problem [6, 7, 8], which may provide optimal switching plans but requires accurate models and becomes computationally prohibitive for large or dynamically changing feeders [9]. Moreover, restoration is inherently sequential and stochastic, with each switching action altering future feasibility [10]. Reinforcement learning (RL) provides an alternative by learning control policies directly through interaction [11], supported by broader advances in deep and graph-based learning for power systems [12]. Early RL efforts used single-agent controllers based on DQN [13] or actor–critic methods [14], but these approaches suffered from biased value estimates [15], inefficient exploration, and limited scalability in feeders with many controllable devices [16].
Multi-agent reinforcement learning (MARL) addresses scalability by assigning local controllers to network subregions. Methods such as MAES [17], MAGDPG [18], Mean-Field RL [19], and QMIX [20] improve coordination but still face key limitations: (i) reliance on value-based updates or fixed exploration rules, which amplify approximation errors as agents learn simultaneously; (ii) feasibility enforcement via action masking or early termination, which disrupts learning and leads to unsafe or infeasible actions; and (iii) overfitting to common operating patterns, reducing performance under unseen conditions.
To overcome these challenges, this paper applies the Heterogeneous-Agent Proximal Policy Optimization (HAPPO) algorithm within a heterogeneous-agent reinforcement learning (HARL) architecture [22] for power distribution system restoration. HAPPO extends PPO to multi-agent settings through sequential per-agent updates guided by a centralized critic, enabling stable and sample-efficient learning among microgrids with differing load levels, generation capacities, and switch counts. These structural differences make agents non-exchangeable and unsuitable for parameter-sharing methods such as MAPPO.
To enable physics-informed, constraint-aware learning, we develop an OpenDSS-based simulation environment [23] that performs full three-phase power-flow analysis after each switching action. Rather than masking infeasible actions [24] or terminating episodes [15], the environment applies a differentiable penalty term proportional to voltage, current, and generation-limit violations, while enforcing a 2400,kW system-level generation cap and local microgrid balance constraints. This soft-constraint design provides dense feedback and allows agents to learn corrective behavior under infeasible conditions.
Experiments on the IEEE 123-bus [25] and IEEE 8500-node [26] feeders show that the proposed HAPPO framework achieves faster convergence, higher restored power, and improved stability compared with established MARL baselines including DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20]. The contributions of this work are: (1) the first application of HAPPO to feeder-level restoration with structurally heterogeneous microgrid agents; (2) a stable and scalable learning framework combining decentralized actors with a centralized advantage critic; (3) a physics-informed OpenDSS environment with continuous differentiable constraint handling; and (4) extensive evaluation on IEEE test feeders demonstrating superior restoration performance and reproducibility.
II Problem Formulation
II-A Network Model and Decision Variables
Let the distribution network be represented by a graph $G=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ and $\mathcal{E}$ denote buses and branches, respectively. Loads lie on $\mathcal{D}\subseteq\mathcal{N}$ and distributed energy resources (DERs) on $\mathcal{G}\subseteq\mathcal{N}$. A subset $\mathcal{S}\subseteq\mathcal{E}$ contains controllable sectionalizing or tie switches, each with binary state $b_{q}\in\{0,1\}$ indicating a closed (energized) or open (de-energized) line. After an outage, restoration proceeds through discrete switching actions that reconfigure the topology and determine the energized buses and restored loads.
II-B Scenario Randomization and Priorities
Each simulation episode introduces variability by randomizing the faulted or unavailable branches, the placement of critical loads, and any optional changes in DER penetration levels.
Each load $k\in\mathcal{D}$ is assigned a priority weight $c_{k}\in\{1,\dots,10\}$, indicating its relative importance in the restoration process. These priority assignments differ between feeders and determine the weighted-load objective defined below.
II-C Restoration Objective
Let $P_{k}$ denote the restored active power at load bus $k$. The system-level restoration objective is to maximize the weighted restored power:
$\max_{\{b_{q}\}}J=\sum_{k\in\mathcal{D}}c_{k}P_{k},$ (1)
where $J$ measures the total service recovery prioritized by load criticality.
II-D Operational Constraints
At each switching configuration, the distribution system must satisfy fundamental electrical and operational limits.
II-D1 Power Balance
The total restored load and network losses must not exceed the available DER generation:
$\sum_{k\in\mathcal{D}}P_{k}+\sum_{\ell\in\mathcal{E}}P^{\text{loss}}_{\ell}\leq\sum_{g\in\mathcal{G}}P_{g}.$ (C1)
II-D2 Voltage Limits
Voltage magnitude at each bus must remain within allowable bounds:
$V_{\min}\leq V_{i}\leq V_{\max},\quad\forall i\in\mathcal{N}.$ (C2)
II-D3 DER Operating Limits
Each DER must operate within its active and reactive power bounds:
$P_{g}^{\min}\leq P_{g}\leq P_{g}^{\max},\quad Q_{g}^{\min}\leq Q_{g}\leq Q_{g}^{\max},\quad\forall g\in\mathcal{G}.$ (C3)
II-D4 Thermal Line Limits
Branch power flows must satisfy thermal constraints:
$P_{ij}^{2}+Q_{ij}^{2}\leq(S^{\max}_{ij})^{2},\quad\forall(i,j)\in\mathcal{E}.$ (C4)
II-D5 Global Generation Capacity
The total DER output across all microgrids cannot exceed the system-wide generation cap:
$\sum_{g\in\mathcal{G}}P_{g}\leq 2400~\text{kW}.$ (C5)
II-D6 Local Microgrid Balance
For each microgrid $m$, local load consumption cannot exceed its local DER generation:
$P^{\text{load}}_{m}\leq P^{\text{gen}}_{m},\quad\forall m.$ (C6)
These constraints define the feasible space of valid switching configurations during restoration.
II-E Structural Heterogeneity and Feeder Partitioning
To reduce the combinatorial complexity of the restoration problem, the feeder is partitioned into $N$ microgrids. Each microgrid includes its own subset of loads and DERs, operates a distinct set of controllable switches $\mathcal{Q}_{i}\subseteq\mathcal{S}$, and exhibits unique load magnitudes and generation capacities, resulting in structurally heterogeneous operating regions. Although all switches share the same action type (open/close/no-op), the number of switches, load profiles, DER availability, and topological location vary significantly across microgrids. This introduces structural heterogeneity, making microgrids non-exchangeable and motivating the heterogeneous-agent formulation adopted in later sections.
II-F Reward Function (for RL-Based Optimization)
When solving (1) via reinforcement learning, a dense reward function is constructed to encourage restoration progress while penalizing constraint violations $r_{t}=\alpha\,\Delta P^{\text{rest}}_{t}-\beta\frac{P^{\text{loss}}_{t}}{P_{\text{gen}}}-\lambda\,\xi_{t},$ where $\Delta P^{\text{rest}}_{t}$ represents the incremental increase in restored power, and $P_{\text{loss}}$ is normalized by the fixed generation cap $P_{\text{gen}}=2400~\text{kW}$. The penalty term $\xi_{t}$ collects the magnitudes of violations of constraints (C1)–(C6), whereas the parameters $\alpha$, $\beta$, and $\lambda$ serve as nonnegative weights that tune the contributions of restoration progress, loss minimization, and constraint enforcement, respectively.
This reward formulation preserves the mathematical objective while enabling learning-based optimization in Section III.
III Proposed HAPPO-Based Methodology
III-A HARL Architecture Overview
The proposed approach is built within a heterogeneous-agent reinforcement learning (HARL) architecture [22], where each microgrid agent maintains an independent actor policy and a shared centralized critic estimates global value functions. This design reflects the physical hierarchy of distribution systems: local agents observe and control only their own microgrids, while the critic leverages system-wide information to provide coordinated learning signals.
Each agent $i$ follows a policy $a_{i,t}=\pi_{\theta_{i}}(o_{i,t}),$ issuing a discrete switching action (open, close, or no-op). The centralized critic $V_{\phi}(s_{t})$ processes the global system state to compute value and advantage estimates that guide all agents.
A defining feature of HAPPO is its sequential per-agent update rule: agents update their policies one at a time, with each update conditioned on the latest policies of previously updated agents. This reduces gradient interference and enhances stability in strongly coupled multi-agent settings such as distribution feeders.
Figure 1: Overview of the proposed HAPPO-based HARL framework for multi-agent power distribution system restoration. Individual microgrid actors issue switching actions based on local observations, while a shared centralized critic computes value and advantage estimates from global states. OpenDSS [23] simulation feedback enables constraint-aware sequential policy updates.
III-B Agent–Environment Interaction
At each simulation step $t$, each agent receives a local observation $o_{i,t}$ (e.g., voltages, currents, DER outputs, and switch states), selects an action $a_{i,t}=\pi_{\theta_{i}}(o_{i,t})$, and the joint action $a_{t}=(a_{1,t},\dots,a_{N,t})$ is executed in the OpenDSS simulator. OpenDSS computes the resulting feeder state, including next global state $s_{t+1}$, updated local observations $\{o_{i,t+1}\}$, restored power and system losses, voltage and current magnitudes, and the shared global reward $r_{t}$.
Each transition contributes to an on-policy trajectory buffer $(o^{i}_{t},a^{i}_{t},r_{t},o^{i}_{t+1},\psi_{t}),$ where $\psi_{t}$ contains shared global signals such as $s_{t}$ or the joint action $a_{t}$. These trajectories support centralized advantage computation and sequential policy optimization.
III-C Advantage Computation via $\lambda$-GAE
To quantify each agent’s temporal contribution to the global return, the framework employs the generalized advantage estimator (GAE):
$\delta_{t}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (2)
$\hat{A}^{i}_{t}$ $=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\,\delta_{t+\ell},$ (3)
where $\gamma$ is the discount factor and $\lambda\in[0,1]$ controls the bias–variance trade-off. The centralized critic provides low-variance, system-consistent value estimates, enabling each agent to compute advantages aligned with the global restoration objective.
III-D Sequential Policy Updates (HAPPO)
HAPPO performs sequential PPO-style policy updates for agents $1$ through $N$. For agent $i$, the clipped surrogate objective is:
$L_{i}(\theta_{i})=\mathbb{E}\!\left[\min\!\left(r_{i,t}\hat{A}^{i}_{t},\,\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)\hat{A}^{i}_{t}\right)\right]-\beta_{\mathrm{ent}}\,\mathbb{E}[H(\pi_{\theta_{i}})],$ (4)
where the probability ratio is $r_{i,t}=\dfrac{\pi_{\theta_{i}}(a_{i,t}\mid o_{i,t})}{\pi^{\text{old}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}.$
The clipping term enforces a trust-region constraint, preventing large, destabilizing policy shifts, while the entropy term $\beta_{\mathrm{ent}}$ encourages adequate exploration. Sequential updates reduce gradient interference and support monotonic policy improvement.
Critic update: The critic parameters $\phi$ are trained by regressing $V_{\phi}(s_{t})$ onto bootstrapped returns:
$L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right],\qquad\hat{R}_{t}=\hat{A}_{t}+V_{\phi}(s_{t}).$ (5)
III-E Training Workflow
Algorithm 1 summarizes the unified HAPPO training procedure used for both IEEE 123-bus and IEEE 8500-node feeders. Each training iteration consists of several steps: agents first interact with OpenDSS for $T$ steps to generate rollouts; the centralized critic then computes temporal-difference errors and $\lambda$-GAE advantages; each agent subsequently updates its actor parameters through sequential clipped PPO-style optimization; the critic is regressed against bootstrapped returns to refine value estimates; and finally, the updated actor and critic networks are deployed for the next episode. This unified workflow enables the method to scale seamlessly between medium- and large-scale feeders, ensuring stable, reproducible learning dynamics.
1
Input: Environment $\mathcal{E}\in\{\text{IEEE-123},\text{IEEE-8500}\}$, configuration $\mathcal{C}$
Output: Trained actor and critic networks for all agents
2
3Initialize: agents, critic $V_{\phi}$, rollout buffer $\mathcal{B}$, and logger;
4 Load hyperparameters: rollout length $T$, learning rates, clipping threshold $\epsilon$, discount $\gamma$, GAE parameter $\lambda$, etc.;
5
6for each iteration $u=1$ to $U_{\max}$ do
7  Reset environment and obtain $(o_{k,0},s_{0})$;
8
9 for $t=0$ to T1T-1 do
10     foreach agent $k$ do
11        Sample action $a_{k,t}\sim\pi_{\theta_{k}}(a|o_{k,t})$;
12
13    Execute joint action $\mathbf{a}_{t}$ in OpenDSS;
14     Observe $(o_{k,t+1},s_{t+1},r_{t},d_{t},\psi_{t})$;
15     Store transition in buffer $\mathcal{B}$;
16
17
18 Compute TD errors and GAE advantages: $\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\quad\hat{A}_{t}=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\delta_{t+\ell}$
19 foreach agent $k$ (sequential update) do
20     Freeze $\pi_{\theta_{j}}$ for all $j\neq k$;
21     for each PPO epoch do
22        Sample minibatches $(o_{k},a_{k},\hat{A},\log\pi_{\text{old}})$;
23        Compute ratio $r$ and clipped surrogate objective;
24        Update actor parameters $\theta_{k}\leftarrow\theta_{k}-\alpha_{\pi}\nabla L_{\text{actor}}$;
25
26
27
28 Compute critic loss: $L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-(\hat{A}_{t}+V_{\phi}(s_{t})))^{2}\right]$ Update critic: $\phi\leftarrow\phi-\alpha_{V}\nabla L_{\text{critic}}$;
29
30 Log training metrics and save checkpoints;
31
32
33return trained $\{\pi_{\theta_{k}}\}_{k=1}^{N}$ and critic $V_{\phi}$;
Algorithm 1 Unified HAPPO Training Workflow for IEEE-123 and IEEE-8500 Feeders
IV Experimental Setup
IV-A Simulation Environment
Experiments are conducted on the IEEE 123-bus [25] and IEEE 8500-node [26] feeders using a custom Python–OpenDSS [23] environment that performs full three-phase AC power-flow calculations after each switching action. The simulator returns voltages, currents, DER outputs, losses, and restored load. Operational constraints, including voltage limits, thermal ratings, DER bounds, and the global generation cap $P_{\text{gen}}=2400\,\text{kW}$, are enforced through differentiable penalty terms rather than action masking or early termination, providing dense feedback near feasibility boundaries.
The IEEE 123-bus feeder (Fig. 2) contains 123 buses, a 3025 kW total load, five DER units with 2400 kW combined capacity, and 26 controllable switches. For multi-agent training, it is divided into five microgrids, each controlled by one agent responsible for local switching. The feeder exhibits realistic unbalanced loading and DER interactions.
Figure 2: IEEE 123-bus distribution feeder used in the experiments. The system includes 26 controllable switches and five DERs (total generation 2400 kW), partitioned into five microgrid regions.
The IEEE 8500-node feeder contains roughly 8500 nodes, a 25 MW total load, 100 controllable switches, and DERs providing up to 2400 kW. It is partitioned into ten microgrids, each managed by a dedicated agent. The large state dimensionality and deep radial topology make it a challenging testbed for evaluating scalable multi-agent coordination with HAPPO.
IV-B Training Configuration
All experiments follow the unified HARL training workflow in Section III (Algorithm 1). Each episode runs for $T$ steps, with decentralized recurrent actors producing discrete switching actions and a centralized critic $V_{\phi}(s_{t})$ computing values and $\lambda$-GAE advantages. Policy learning uses PPO-style updates with clipping threshold $\epsilon$, entropy coefficient $\beta_{\text{ent}}$, and sequential per-agent optimization. Rewards include incremental restored power, normalized loss penalties, and continuous constraint-violation terms. A fixed generation cap of $P_{\text{gen}}=2400\,\text{kW}$ is applied to both feeders. Hyperparameters ($\gamma$, $\lambda$, learning rates, batch sizes, etc.) remain identical across systems unless scale adjustments are required.
IV-C Hardware and Implementation Details
Training is performed on a workstation with an NVIDIA RTX 6000 Ada GPU (48 GB), AMD Threadripper PRO CPU, and 256 GB RAM. OpenDSS runs on the CPU, while PyTorch handles actor–critic updates on the GPU. The HARL framework supports decentralized actors, a centralized critic, sequential PPO updates, and synchronized rollout collection.
IV-D Evaluation Metrics
Performance is evaluated using cumulative reward (weighted restoration benefit), restored active power (relative to the 2400 kW cap), convergence stability across seeds, inference latency per switching action, and total training time. This unified protocol enables consistent comparison across feeders and baseline RL/MARL methods.
V Numerical Results
This section evaluates the proposed HAPPO-based restoration framework on the IEEE 123-bus and IEEE 8500-node systems using the unified HARL training pipeline described in Sections IIIIV. Performance is reported in terms of (i) cumulative reward, which reflects the weighted restoration benefit accumulated during each episode, and (ii) restored active power, which directly measures physical restoration effectiveness relative to the system-wide generation cap $P_{\text{gen}}=2400\,\text{kW}$. All results are averaged over multiple independent random seeds.
V-A Training Performance on IEEE 123-Bus and IEEE 8500-Node Systems
Figures 3 and 4 show the evolution of cumulative reward and restored active power during training for the two feeders. Each curve corresponds to an independent seed initialized with randomized outage locations, critical-load placements, and DER conditions.
On both feeders, agents select among three discrete switching actions (open, close, or no-op) and receive a shared reward that incorporates incremental restored power, normalized loss penalties, and differentiable constraint-violation terms. Unlike masking-based MARL methods, the proposed penalty-driven reward ensures uninterrupted learning near feasibility boundaries and supports smooth gradient propagation.
In the IEEE 123-bus feeder 5 agents coordinate 26 switches. Across five independent training runs, the cumulative reward increases steadily and converges to $(3.6\pm 2.9)\times 10^{6},$ while the restored power stabilizes at $2294\pm 52~\text{kW}\quad(95.6\%\pm 2.2\%\text{ of }P_{\text{gen}}).$ As shown in Fig. 3, performance across seeds is tightly clustered, indicating reliable convergence and stable sequential actor updates.
In the IEEE 8500-node feeder, 10 agents each govern ten randomly assigned switches across 1,177 loads. Despite the dramatically larger state–action space and deeper network topology, HAPPO maintains monotonic improvement. The cumulative reward converges to $(2.9\pm 3.3)\times 10^{6},$ and restored power reaches $2308\pm 55~\text{kW}\quad(\!96.2\%\pm 2.3\%\text{ of }P_{\text{gen}}),$ as shown in Fig. 4. These results highlight the scalability of sequential trust-region updates, centralized advantage estimation, and penalty-based reward shaping.
Figure 3: Training performance of HAPPO on the IEEE 123-bus feeder across four random seeds. Left: cumulative reward. Right: restored active power.
Figure 4: Training performance of HAPPO on the IEEE 8500-node feeder across five random seeds. Left: cumulative reward. Right: restored active power.
TABLE I: Performance comparison of HAPPO with state-of-the-art RL/MARL methods on IEEE 123-bus and IEEE 8500-node feeders (generation cap $P_{\text{gen}}=2400\,\text{kW}$)
Method IEEE 123 IEEE 8500
Cum. Reward Restored (%) Train Time (min) Inference (ms) Cum. Reward Restored (%) Train Time (min) Inference (ms)
DQN [13] 28.38 $\pm$ 0.14 73.90 $\pm$ 0.46 34.7 4.2 65.58 $\pm$ 0.25 69.50 $\pm$ 0.28 249.3 6.8
PPO [27] 29.01 $\pm$ 0.15 74.85 $\pm$ 0.47 44.8 7.1 67.94 $\pm$ 0.35 71.05 $\pm$ 0.27 255.7 9.8
MAES [17] 29.78 $\pm$ 0.19 76.14 $\pm$ 0.40 91.8 25.4 69.82 $\pm$ 0.32 72.81 $\pm$ 0.23 307.0 38.7
MAGDPG [18] 30.08 $\pm$ 0.17 77.00 $\pm$ 0.42 118.2 29.5 70.56 $\pm$ 0.24 73.44 $\pm$ 0.32 321.5 45.6
MADQN [24] 30.58 $\pm$ 0.22 77.80 $\pm$ 0.59 98.8 12.6 71.37 $\pm$ 0.34 74.02 $\pm$ 0.39 279.2 19.0
Mean-Field RL [19] 31.56 $\pm$ 0.25 80.26 $\pm$ 0.52 108.3 27.3 73.46 $\pm$ 0.33 77.40 $\pm$ 0.38 303.9 42.6
QMIX [20] 33.38 $\pm$ 0.25 83.25 $\pm$ 0.45 96.4 15.8 76.62 $\pm$ 0.34 80.10 $\pm$ 0.30 291.3 23.5
[…]
V-B Multi-Seed Convergence and Reproducibility
To assess robustness under stochastic outage conditions, each experiment is repeated across multiple random seeds. Across both feeders, the trajectories in Figs. 3 and 4 exhibit consistent monotonic convergence in both metrics.
On the IEEE 123-bus feeder, restored power converges within $2230\text{--}2367~\text{kW},$ corresponding to 93–99% of the generation cap. On the IEEE 8500-node feeder, restored power stabilizes near $2224\text{--}2347~\text{kW},$ representing 93–98% of the available capacity.
These results confirm that the proposed HARL-based HAPPO framework provides highly reproducible performance due to (i) sequential per-agent trust-region updates, (ii) low-variance centralized $\lambda$-GAE advantage estimation, and (iii) penalty-driven constraint handling that avoids disruptive early terminations.
V-C Comparative Evaluation with Baseline Methods
HAPPO is compared with state-of-the-art RL/MARL baselines, including DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20]. All methods are trained under identical outage scenarios, hyperparameters, and generation cap $P_{\text{gen}}=2400\,\text{kW}$. Table I summarizes quantitative performance.
HAPPO achieves the highest restored power and cumulative reward on both feeders. Its sequential per-agent trust-region updates and centralized advantage estimation enable stable policy optimization even in the presence of strong physical coupling and high-dimensional state spaces, outperforming value-based and value-decomposition methods that struggle under these conditions.
V-D Discussion and Interpretation
The results demonstrate four key findings. (1) HAPPO exhibits strong scalability, extending from a five-agent IEEE 123 system to a ten-agent IEEE 8500 system with minimal degradation in convergence smoothness. (2) The differentiable penalty-based reward enables continuous guidance near operational limits, avoiding the brittleness typically observed in masking-based approaches and enhancing constraint awareness. (3) Multi-seed experiments show consistently low variance, confirming the reproducibility of sequential PPO-style updates. (4) The OpenDSS-based simulation environment ensures physical fidelity by enforcing operational limits and producing policies that correspond to feasible feeder configurations in realistic distribution systems.
These findings confirm that the HARL-based HAPPO framework is well suited for large-scale, constraint-driven distribution system restoration.
VI Conclusion
This paper presented a HAPPO approach for distribution system restoration within a HARL framework using an OpenDSS-based multi-agent environment. Sequential trust-region updates combined with centralized advantage estimation enable stable, sample-efficient coordination among structurally heterogeneous microgrid agents, avoiding the instability issues common in conventional MARL. A continuous penalty-based reward replaces masking or early termination, yielding smooth gradients and naturally constraint-aware switching behavior.
Experiments on the IEEE 123-bus and IEEE 8500-node feeders show that HAPPO consistently outperforms PPO, QMIX, Mean-Field RL, and other baselines. Under the 2400 kW generation cap, it restores roughly 92% and 96% of available load on the two systems, respectively, with strong multi-seed reproducibility and low inference latency. The OpenDSS integration ensures physical validity and supports direct transition to operational settings.
Overall, the results demonstrate that heterogeneous-agent policy optimization provides a scalable, stable, and realistic solution for resilient distribution system restoration. Future work will incorporate model-based elements, communication constraints, partial observability, and dynamic DER participation to further strengthen real-time applicability.
Acknowledgment
This research was supported in part by the National Science Foundation under grant ECCS-2223628.

### v5

@@
 Abstract
-Restoring power distribution systems (PDS) after large-scale outages requires sequential switching operations that reconfigure feeder topology and coordinate distributed energy resources (DERs) under nonlinear constraints such as power balance, voltage limits, and thermal ratings.
-These challenges make conventional optimization and value-based RL approaches computationally inefficient and difficult to scale.
-This paper applies a Heterogeneous-Agent Reinforcement Learning (HARL) framework, instantiated through Heterogeneous-Agent Proximal Policy Optimization (HAPPO), to enable coordinated restoration across interconnected microgrids.
-Each agent controls a distinct microgrid with different loads, DER capacities, and switch counts, introducing practical structural heterogeneity.
-Decentralized actor policies are trained with a centralized critic to compute advantage values for stable on-policy updates.
-A physics-informed OpenDSS environment provides full power-flow feedback and enforces operational limits via differentiable penalty signals rather than invalid action masking.
-The total DER generation is capped at 2400 kW, and each microgrid must satisfy local supply–demand feasibility.
-Experiments on the IEEE 123-bus and IEEE 8500-node systems show that HAPPO achieves faster convergence, higher restored power, and smoother multi-seed training than DQN, PPO, MAES, MAGDPG, MADQN, Mean-Field RL, and QMIX.
-Results demonstrate that incorporating microgrid-level heterogeneity within the HARL framework yields a scalable, stable, and constraint-aware solution for complex PDS restoration.
+Restoring power distribution systems (PDSs) after large-scale outages requires sequential switching actions that reconfigure feeder topology and coordinate distributed energy resources (DERs) under nonlinear constraints, including power balance, voltage limits, and thermal ratings.
+These challenges limit the scalability of conventional optimization and value-based reinforcement learning (RL) approaches.
+This paper applies a Heterogeneous-Agent Reinforcement Learning (HARL) framework via Heterogeneous-Agent Proximal Policy Optimization (HAPPO) to enable coordinated restoration across interconnected microgrids.
+Each agent controls a distinct microgrid with different loads, DER capacities, and switch counts.
+Decentralized actors are trained with a centralized critic for stable on-policy learning, while a physics-informed OpenDSS environment enforces electrical feasibility.
+Experiments on IEEE 123-bus and 8500-node feeders show HAPPO outperforms PPO, QMIX, Mean-Field RL, and other baselines in restored power, convergence stability, and multi-seed reproducibility.
+Under a 2400 kW generation cap, the framework restores over 95% of available load on both systems with low-latency execution, supporting practical real-time PDS restoration.
 I Introduction
@@
 Distributed energy resources (DERs) such as photovoltaics, microturbines, and storage can accelerate local recovery [2], yet their limited capacity and the voltage, thermal, and topological constraints of distribution feeders complicate safe and efficient restoration [3].
-Recent advancements in machine learning for critical infrastructure highlight the growing role of data-driven recovery, including a transformer–reinforcement learning framework for large-scale repair sequencing that enables faster and more resilient system restoration [4].
-Machine learning has also advanced modeling in other safety-critical physical domains, including high-fidelity ground-motion prediction using ensemble techniques [5].
-Conventional restoration is typically formulated as a mixed-integer nonlinear optimization problem [6, 7, 8], which may provide optimal switching plans but requires accurate models and becomes computationally prohibitive for large or dynamically changing feeders [9].
+Recent advancements in machine learning for critical infrastructure highlight the growing role of data-driven recovery, including transformer–reinforcement learning frameworks for large-scale repair sequencing [4] and ensemble techniques for high-fidelity ground-motion prediction [5].
+Conventional restoration is typically formulated as a mixed-integer nonlinear optimization problem [6, 7, 8], which may provide optimal switching plans but becomes computationally prohibitive for large or dynamically changing feeders [9].
 Moreover, restoration is inherently sequential and stochastic, with each switching action altering future feasibility [10].
 Reinforcement learning (RL) provides an alternative by learning control policies directly through interaction [11], supported by broader advances in deep and graph-based learning for power systems [12].
-Early RL efforts used single-agent controllers based on DQN [13] or actor–critic methods [14], but these approaches suffered from biased value estimates [15], inefficient exploration, and limited scalability in feeders with many controllable devices [16].
+Early RL efforts used single-agent controllers based on DQN [13] or actor–critic methods [14], but suffered from biased value estimates [15], inefficient exploration, and limited scalability [16].
 Multi-agent reinforcement learning (MARL) addresses scalability by assigning local controllers to network subregions.
-Methods such as MAES [17], MAGDPG [18], Mean-Field RL [19], and QMIX [20] improve coordination but still face key limitations: (i) reliance on value-based updates or fixed exploration rules, which amplify approximation errors as agents learn simultaneously; (ii) feasibility enforcement via action masking or early termination, which disrupts learning and leads to unsafe or infeasible actions; and (iii) overfitting to common operating patterns, reducing performance under unseen conditions.
-To overcome these challenges, this paper applies the Heterogeneous-Agent Proximal Policy Optimization (HAPPO) algorithm within a heterogeneous-agent reinforcement learning (HARL) architecture [22] for power distribution system restoration.
-HAPPO extends PPO to multi-agent settings through sequential per-agent updates guided by a centralized critic, enabling stable and sample-efficient learning among microgrids with differing load levels, generation capacities, and switch counts.
+Methods such as MAES [17], MAGDPG [18], Mean-Field RL [19], and QMIX [20] improve coordination but face key limitations: (i) reliance on value-based updates or fixed exploration rules, amplifying approximation errors; (ii) feasibility enforcement via action masking or early termination, disrupting learning; and (iii) overfitting to common operating patterns.
+To overcome these challenges, this paper applies Heterogeneous-Agent Proximal Policy Optimization (HAPPO) within a heterogeneous-agent reinforcement learning (HARL) architecture [22] for power distribution system restoration.
+HAPPO extends PPO to multi-agent settings through sequential per-agent updates guided by a centralized critic, enabling stable learning among microgrids with differing load levels, generation capacities, and switch counts.
 These structural differences make agents non-exchangeable and unsuitable for parameter-sharing methods such as MAPPO.
-To enable physics-informed, constraint-aware learning, we develop an OpenDSS-based simulation environment [23] that performs full three-phase power-flow analysis after each switching action.
-Rather than masking infeasible actions [24] or terminating episodes [15], the environment applies a differentiable penalty term proportional to voltage, current, and generation-limit violations, while enforcing a 2400,kW system-level generation cap and local microgrid balance constraints.
-This soft-constraint design provides dense feedback and allows agents to learn corrective behavior under infeasible conditions.
-Experiments on the IEEE 123-bus [25] and IEEE 8500-node [26] feeders show that the proposed HAPPO framework achieves faster convergence, higher restored power, and improved stability compared with established MARL baselines including DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20].
-The contributions of this work are: (1) the first application of HAPPO to feeder-level restoration with structurally heterogeneous microgrid agents; (2) a stable and scalable learning framework combining decentralized actors with a centralized advantage critic; (3) a physics-informed OpenDSS environment with continuous differentiable constraint handling; and (4) extensive evaluation on IEEE test feeders demonstrating superior restoration performance and reproducibility.
+We develop an OpenDSS-based simulation environment [23] performing full three-phase power-flow analysis after each switching action.
+Rather than masking infeasible actions [24] or terminating episodes [15], the environment applies differentiable penalty terms proportional to voltage, current, and generation-limit violations, while enforcing a system-level generation cap and local microgrid balance constraints.
+This soft-constraint design provides dense feedback and allows agents to learn corrective behavior.
+Experiments on IEEE 123-bus [25] and 8500-node [26] feeders show the proposed HAPPO framework achieves faster convergence, higher restored power, and improved stability compared with DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20].
+The contributions are: (1) first application of HAPPO to feeder-level restoration with structurally heterogeneous microgrid agents; (2) a stable and scalable learning framework combining decentralized actors with a centralized advantage critic.
 II Problem Formulation
 II-A Network Model and Decision Variables
-Let the distribution network be represented by a graph $G=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ and $\mathcal{E}$ denote buses and branches, respectively.
+Let the distribution network be represented by a graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ denotes the set of buses and $\mathcal{E}$ denotes the set of branches.
 Loads lie on $\mathcal{D}\subseteq\mathcal{N}$ and distributed energy resources (DERs) on $\mathcal{G}\subseteq\mathcal{N}$.
-A subset $\mathcal{S}\subseteq\mathcal{E}$ contains controllable sectionalizing or tie switches, each with binary state $b_{q}\in\{0,1\}$ indicating a closed (energized) or open (de-energized) line.
+A subset $\mathcal{S}\subseteq\mathcal{E}$ contains controllable sectionalizing or tie switches.
+Each switch $q\in\mathcal{S}$ has a binary state $b_{q}\in\{0,1\}$ indicating a closed (energized) or open (de-energized) line.
 After an outage, restoration proceeds through discrete switching actions that reconfigure the topology and determine the energized buses and restored loads.
@@
 These priority assignments differ between feeders and determine the weighted-load objective defined below.
-II-C Restoration Objective
+II-C Restoration Objective and Operational Constraints
 Let $P_{k}$ denote the restored active power at load bus $k$. The system-level restoration objective is to maximize the weighted restored power:
@@
 where $J$ measures the total service recovery prioritized by load criticality.
-II-D Operational Constraints
 At each switching configuration, the distribution system must satisfy fundamental electrical and operational limits.
-II-D1 Power Balance
-The total restored load and network losses must not exceed the available DER generation:
-$\sum_{k\in\mathcal{D}}P_{k}+\sum_{\ell\in\mathcal{E}}P^{\text{loss}}_{\ell}\leq\sum_{g\in\mathcal{G}}P_{g}.$ (C1)
-II-D2 Voltage Limits
+These constraints define the feasible space of valid switching configurations during restoration.
+II-C1 Power Balance
+The total restored load and network losses must not exceed the available DER generation. Here, the inequality permits unused generation capacity during partial restoration:
+$\sum_{k\in\mathcal{D}}P_{k}+\sum_{e\in\mathcal{E}}P^{\text{loss}}_{e}\leq\sum_{g\in\mathcal{G}}P_{g},$ (2)
+where $P_{k}$ is the active power consumed by load $k$, $P^{\text{loss}}_{e}$ is the power loss on branch $e$, and $P_{g}$ is the active power generated by DER unit $g$.
+II-C2 Voltage Limits
 Voltage magnitude at each bus must remain within allowable bounds:
-$V_{\min}\leq V_{i}\leq V_{\max},\quad\forall i\in\mathcal{N}.$ (C2)
-II-D3 DER Operating Limits
+$V_{\min}\leq V_{k}\leq V_{\max},\quad\forall k\in\mathcal{N},$ (3)
+where $V_{k}$ is the voltage magnitude at bus $k$.
+II-C3 DER Operating Limits
 Each DER must operate within its active and reactive power bounds:
-$P_{g}^{\min}\leq P_{g}\leq P_{g}^{\max},\quad Q_{g}^{\min}\leq Q_{g}\leq Q_{g}^{\max},\quad\forall g\in\mathcal{G}.$ (C3)
-II-D4 Thermal Line Limits
+$P_{g}^{\min}\leq P_{g}\leq P_{g}^{\max},\quad Q_{g}^{\min}\leq Q_{g}\leq Q_{g}^{\max},\quad\forall g\in\mathcal{G},$
+where $P_{g}$ and $Q_{g}$ are the active and reactive power outputs of DER unit $g$, respectively.
+II-C4 Thermal Line Limits
 Branch power flows must satisfy thermal constraints:
-$P_{ij}^{2}+Q_{ij}^{2}\leq(S^{\max}_{ij})^{2},\quad\forall(i,j)\in\mathcal{E}.$ (C4)
-II-D5 Global Generation Capacity
+$P_{e}^{2}+Q_{e}^{2}\leq(S^{\max}_{e})^{2},\quad\forall e\in\mathcal{E},$ (5)
+where $P_{e}$ and $Q_{e}$ are the active and reactive power flows on branch $e$, and $S^{\max}_{e}$ is the thermal capacity of branch $e$.
+II-C5 Global Generation Capacity
 The total DER output across all microgrids cannot exceed the system-wide generation cap:
-$\sum_{g\in\mathcal{G}}P_{g}\leq 2400~\text{kW}.$ (C5)
-II-D6 Local Microgrid Balance
-For each microgrid $m$, local load consumption cannot exceed its local DER generation:
-$P^{\text{load}}_{m}\leq P^{\text{gen}}_{m},\quad\forall m.$ (C6)
-These constraints define the feasible space of valid switching configurations during restoration.
-II-E Structural Heterogeneity and Feeder Partitioning
+$\sum_{g\in\mathcal{G}}P_{g}\leq 2400~\text{kW}.$ (6)
+II-C6 Local Microgrid Balance
+For each microgrid $i$, local load consumption cannot exceed its local DER generation:
+$P^{\text{load}}_{i}\leq P^{\text{gen}}_{i},\quad\forall i.$ (7)
+II-D Structural Heterogeneity and Feeder Partitioning
 To reduce the combinatorial complexity of the restoration problem, the feeder is partitioned into $N$ microgrids.
-Each microgrid includes its own subset of loads and DERs, operates a distinct set of controllable switches $\mathcal{Q}_{i}\subseteq\mathcal{S}$, and exhibits unique load magnitudes and generation capacities, resulting in structurally heterogeneous operating regions.
-Although all switches share the same action type (open/close/no-op), the number of switches, load profiles, DER availability, and topological location vary significantly across microgrids.
-This introduces structural heterogeneity, making microgrids non-exchangeable and motivating the heterogeneous-agent formulation adopted in later sections.
-II-F Reward Function (for RL-Based Optimization)
-When solving (1) via reinforcement learning, a dense reward function is constructed to encourage restoration progress while penalizing constraint violations $r_{t}=\alpha\,\Delta P^{\text{rest}}_{t}-\beta\frac{P^{\text{loss}}_{t}}{P_{\text{gen}}}-\lambda\,\xi_{t},$ where $\Delta P^{\text{rest}}_{t}$ represents the incremental increase in restored power, and $P_{\text{loss}}$ is normalized by the fixed generation cap $P_{\text{gen}}=2400~\text{kW}$.
-The penalty term $\xi_{t}$ collects the magnitudes of violations of constraints (C1)–(C6), whereas the parameters $\alpha$, $\beta$, and $\lambda$ serve as nonnegative weights that tune the contributions of restoration progress, loss minimization, and constraint enforcement, respectively.
+Each microgrid $i$ includes its own subset of loads and DERs, operates a distinct set of controllable switches $\mathcal{Q}_{i}\subseteq\mathcal{S}$, and exhibits unique load magnitudes and generation capacities, resulting in structurally heterogeneous operating regions.
+While each agent controls switches with identical action types (open/close/no-op), microgrids differ substantially in the number of switches they control, their load profiles, DER availability, and topological structure.
+This structural heterogeneity makes microgrids non-exchangeable and motivates the heterogeneous-agent formulation adopted in later sections.
+Microgrids are defined following standard feeder partitioning practices based on topology and DER locality, consistent with prior restoration benchmarks to ensure fair comparison.
+This choice reflects physical connectivity and operational coupling, and alternative reasonable partitions yielded similar convergence behavior, indicating low sensitivity to minor variations.
+II-E Reward Function
+When solving (1) via reinforcement learning, a dense reward function is constructed to encourage restoration progress while penalizing constraint violations.
+Let $\Delta P_{t}^{\text{rest}}=\sum_{k\in\mathcal{D}}c_{k}(P_{k,t}-P_{k,t-1})$ denote the weighted incremental increase in restored power at step $t$, where $P_{k,t}$ is the power restored to load $k$ at time $t$.
+Then $r_{t}=\alpha\,\Delta P^{\text{rest}}_{t}-\beta\frac{P^{\text{loss}}_{t}}{P_{\text{gen}}}-\lambda\,\xi_{t},$ where $\Delta P^{\text{rest}}_{t}$ represents the incremental increase in restored power, and $P^{\text{loss}}$ is normalized by the fixed generation cap $P_{\text{gen}}=2400~\text{kW}$.
+The penalty term $\xi_{t}$ collects the magnitudes of violations of constraints (2)–(7), whereas the parameters $\alpha$, $\beta$, and $\lambda$ serve as nonnegative weights that tune the contributions of restoration progress, loss minimization, and constraint enforcement, respectively.
 This reward formulation preserves the mathematical objective while enabling learning-based optimization in Section III.
@@
 This design reflects the physical hierarchy of distribution systems: local agents observe and control only their own microgrids, while the critic leverages system-wide information to provide coordinated learning signals.
-Each agent $i$ follows a policy $a_{i,t}=\pi_{\theta_{i}}(o_{i,t}),$ issuing a discrete switching action (open, close, or no-op).
+Each agent $i$ follows a policy $a_{i,t}\sim\pi_{\theta_{i}}(o_{i,t}),$ issuing a discrete switching action (open, close, or no-op).
 The centralized critic $V_{\phi}(s_{t})$ processes the global system state to compute value and advantage estimates that guide all agents.
@@
 III-B Agent–Environment Interaction
-At each simulation step $t$, each agent receives a local observation $o_{i,t}$ (e.g., voltages, currents, DER outputs, and switch states), selects an action $a_{i,t}=\pi_{\theta_{i}}(o_{i,t})$, and the joint action $a_{t}=(a_{1,t},\dots,a_{N,t})$ is executed in the OpenDSS simulator.
+At each simulation step $t$, each agent $i$ receives a local observation $o_{i,t}$ (e.g., voltages, currents, DER outputs, and switch states), selects an action $a_{i,t}\sim\pi_{\theta_{i}}(o_{i,t})$, and the joint action $a_{t}=(a^{1}_{t},\ldots,a^{N}_{t})$ is executed in the OpenDSS simulator.
 OpenDSS computes the resulting feeder state, including next global state $s_{t+1}$, updated local observations $\{o_{i,t+1}\}$, restored power and system losses, voltage and current magnitudes, and the shared global reward $r_{t}$.
-Each transition contributes to an on-policy trajectory buffer $(o^{i}_{t},a^{i}_{t},r_{t},o^{i}_{t+1},\psi_{t}),$ where $\psi_{t}$ contains shared global signals such as $s_{t}$ or the joint action $a_{t}$.
+Each transition contributes to an on-policy trajectory buffer $(o^{i}_{t},a^{i}_{t},r_{t},o^{i}_{t+1},\psi_{t}),$ where $\psi_{t}$ contains auxiliary information such as the joint action $\mathbf{a}_{t}=(a_{1,t},\ldots,a_{N,t})$ and other global signals used for centralized training.
 These trajectories support centralized advantage computation and sequential policy optimization.
@@
 To quantify each agent’s temporal contribution to the global return, the framework employs the generalized advantage estimator (GAE):
-$\delta_{t}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (2)
-$\hat{A}^{i}_{t}$ $=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\,\delta_{t+\ell},$ (3)
-where $\gamma$ is the discount factor and $\lambda\in[0,1]$ controls the bias–variance trade-off.
+$\delta_{t}^{i}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (8)
+$\hat{A}^{i}_{t}$ $=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\,\delta_{t+\ell}^{i},$ (9)
+where $\gamma$ is the discount factor and $\lambda\in[0,1]$ controls the bias–variance trade-off, and $\ell$ is the time-step offset.
+Note that while the TD error uses the global reward $r_{t}$ and centralized value function $V_{\phi}(s_{t})$, advantages are computed separately for each agent $i$ to support sequential policy updates.
 The centralized critic provides low-variance, system-consistent value estimates, enabling each agent to compute advantages aligned with the global restoration objective.
@@
 HAPPO performs sequential PPO-style policy updates for agents $1$ through $N$. For agent $i$, the clipped surrogate objective is:
-$L_{i}(\theta_{i})=\mathbb{E}\!\left[\min\!\left(r_{i,t}\hat{A}^{i}_{t},\,\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)\hat{A}^{i}_{t}\right)\right]-\beta_{\mathrm{ent}}\,\mathbb{E}[H(\pi_{\theta_{i}})],$ (4)
-where the probability ratio is $r_{i,t}=\dfrac{\pi_{\theta_{i}}(a_{i,t}\mid o_{i,t})}{\pi^{\text{old}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}.$
+$L_{i}(\theta_{i})=\mathbb{E}\!\left[\min\!\left(\rho_{i,t}\hat{A}^{i}_{t},\,\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)\hat{A}^{i}_{t}\right)\right]-\beta_{\mathrm{ent}}\,\mathbb{E}[H(\pi_{\theta_{i}})],$
+where the probability ratio is $\rho_{i,t}=\dfrac{\pi^{\text{new}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}{\pi^{\text{old}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}.$ The entropy regularization term $H(\pi_{\theta_{i}})$ encourages exploration, with $\beta_{\mathrm{ent}}$ controlling the strength of the entropy bonus.
 The clipping term enforces a trust-region constraint, preventing large, destabilizing policy shifts, while the entropy term $\beta_{\mathrm{ent}}$ encourages adequate exploration.
 Sequential updates reduce gradient interference and support monotonic policy improvement.
-Critic update: The critic parameters $\phi$ are trained by regressing $V_{\phi}(s_{t})$ onto bootstrapped returns:
-$L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right],\qquad\hat{R}_{t}=\hat{A}_{t}+V_{\phi}(s_{t}).$ (5)
+The critic parameters $\phi$ are trained by regressing $V_{\phi}(s_{t})$ onto bootstrapped returns:
+$L_{\text{critic}}=\mathbb{E}\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right],\quad\hat{R}_{t}=\sum_{\ell=0}^{T-t-1}\gamma^{\ell}r_{t+\ell}+\gamma^{T-t}V_{\phi}(s_{T}).$
 III-E Training Workflow
-Algorithm 1 summarizes the unified HAPPO training procedure used for both IEEE 123-bus and IEEE 8500-node feeders.
-Each training iteration consists of several steps: agents first interact with OpenDSS for $T$ steps to generate rollouts; the centralized critic then computes temporal-difference errors and $\lambda$-GAE advantages; each agent subsequently updates its actor parameters through sequential clipped PPO-style optimization; the critic is regressed against bootstrapped returns to refine value estimates; and finally, the updated actor and critic networks are deployed for the next episode.
-This unified workflow enables the method to scale seamlessly between medium- and large-scale feeders, ensuring stable, reproducible learning dynamics.
-1
+Algorithm 1 summarizes the unified HAPPO training procedure for both IEEE 123-bus and 8500-node feeders.
+Each iteration involves: agents interact with OpenDSS for $T$ steps to generate rollouts; the centralized critic computes temporal-difference errors and $\lambda$-GAE advantages; agents sequentially update actor parameters via clipped PPO-style optimization; the critic regresses against bootstrapped returns to refine value estimates; and updated networks deploy for the next episode.
+This workflow enables seamless scaling between medium- and large-scale feeders, ensuring stable, reproducible learning dynamics.
 Input: Environment $\mathcal{E}\in\{\text{IEEE-123},\text{IEEE-8500}\}$, configuration $\mathcal{C}$
 Output: Trained actor and critic networks for all agents
-2
-3Initialize: agents, critic $V_{\phi}$, rollout buffer $\mathcal{B}$, and logger;
-4 Load hyperparameters: rollout length $T$, learning rates, clipping threshold $\epsilon$, discount $\gamma$, GAE parameter $\lambda$, etc.;
-5
-6for each iteration $u=1$ to $U_{\max}$ do
-7  Reset environment and obtain $(o_{k,0},s_{0})$;
-8
+1
+21exInitialize: agents, critic $V_{\phi}$, rollout buffer $\mathcal{B}$, and logger;
+3 Load hyperparameters: rollout length $T$, learning rates, clipping threshold $\epsilon$, discount $\gamma$, GAE parameter $\lambda$, etc., and checkpoint interval $S$ (save period);
+4
+51exfor each iteration $u=1$ to $U_{\max}$ do
+6  Reset environment and obtain $(o_{i,0},s_{0})$;
 7  for $t=0$ to T1T-1 do
-10     foreach agent $k$ do
-11        Sample action $a_{k,t}\sim\pi_{\theta_{k}}(a|o_{k,t})$;
-12
+8     foreach agent $i$ do
+9        Sample action $a_{i,t}\sim\pi_{\theta_{i}}(a|o_{i,t})$;
+10
 11    Execute joint action $\mathbf{a}_{t}$ in OpenDSS;
-14     Observe $(o_{k,t+1},s_{t+1},r_{t},d_{t},\psi_{t})$;
+12     Observe $(o_{i,t+1},s_{t+1},r_{t},d_{t},\psi_{t})$;
 13     Store transition in buffer $\mathcal{B}$;
+14
+15 Compute TD errors and GAE advantages for each agent: $\delta_{t}^{i}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\quad\hat{A}_{t}^{i}=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\delta_{t+\ell}^{i}$;
 16
-17
-18 Compute TD errors and GAE advantages: $\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\quad\hat{A}_{t}=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\delta_{t+\ell}$
-19 foreach agent $k$ (sequential update) do
-20     Freeze $\pi_{\theta_{j}}$ for all $j\neq k$;
+17 foreach agent $i$ (sequential update) do
+18     Freeze $\pi_{\theta_{j}}$ for all $j\neq i$;
 19     for each PPO epoch do
-22        Sample minibatches $(o_{k},a_{k},\hat{A},\log\pi_{\text{old}})$;
-23        Compute ratio $r$ and clipped surrogate objective;
-24        Update actor parameters $\theta_{k}\leftarrow\theta_{k}-\alpha_{\pi}\nabla L_{\text{actor}}$;
-25
-26
-27
-28 Compute critic loss: $L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-(\hat{A}_{t}+V_{\phi}(s_{t})))^{2}\right]$ Update critic: $\phi\leftarrow\phi-\alpha_{V}\nabla L_{\text{critic}}$;
-29
-30 Log training metrics and save checkpoints;
-31
+20        Sample minibatches $(o_{i},a_{i},\hat{A},\log\pi_{\text{old}})$;
+21        Compute probability ratio $\rho$ and clipped surrogate objective;
+22        Update actor parameters $\theta_{i}\leftarrow\theta_{i}-\alpha_{\pi}\nabla L_{\text{actor}}$;
+23
+24
+25 Compute critic loss: $L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right]$,
+26  where $\hat{R}_{t}=\sum_{\ell=0}^{T-t-1}\gamma^{\ell}r_{t+\ell}+\gamma^{T-t}V_{\phi}(s_{T})$;;
+27  Update critic: $\phi\leftarrow\phi-\alpha_{V}\nabla L_{\text{critic}}$;
+28
+29 Log training metrics at iteration $u$ (iter);
+30  if $u\bmod S=0$ then
+31     Save model checkpoint (ckpt);
 32
-33return trained $\{\pi_{\theta_{k}}\}_{k=1}^{N}$ and critic $V_{\phi}$;
+33
+34return trained $\{\pi_{\theta_{i}}\}_{i=1}^{N}$ and critic $V_{\phi}$;
 Algorithm 1 Unified HAPPO Training Workflow for IEEE-123 and IEEE-8500 Feeders
@@
 The IEEE 123-bus feeder (Fig.
-2) contains 123 buses, a 3025 kW total load, five DER units with 2400 kW combined capacity, and 26 controllable switches.
-For multi-agent training, it is divided into five microgrids, each controlled by one agent responsible for local switching.
-The feeder exhibits realistic unbalanced loading and DER interactions.
+2) has 123 buses, a 3025 kW load, five DERs totaling 2400 kW, and 26 controllable switches, partitioned into five agent-controlled microgrids.
+The IEEE 8500-node feeder contains about 8500 nodes, a 25 MW load, 100 switches, and 2400 kW of DER capacity, divided into ten microgrids.
+Its scale and deep radial structure make it a demanding testbed for HAPPO scalability.
 Figure 2: IEEE 123-bus distribution feeder used in the experiments. The system includes 26 controllable switches and five DERs (total generation 2400 kW), partitioned into five microgrid regions.
-The IEEE 8500-node feeder contains roughly 8500 nodes, a 25 MW total load, 100 controllable switches, and DERs providing up to 2400 kW.
-It is partitioned into ten microgrids, each managed by a dedicated agent.
-The large state dimensionality and deep radial topology make it a challenging testbed for evaluating scalable multi-agent coordination with HAPPO.
 IV-B Training Configuration
-All experiments follow the unified HARL training workflow in Section III (Algorithm 1).
-Each episode runs for $T$ steps, with decentralized recurrent actors producing discrete switching actions and a centralized critic $V_{\phi}(s_{t})$ computing values and $\lambda$-GAE advantages.
-Policy learning uses PPO-style updates with clipping threshold $\epsilon$, entropy coefficient $\beta_{\text{ent}}$, and sequential per-agent optimization.
-Rewards include incremental restored power, normalized loss penalties, and continuous constraint-violation terms.
-A fixed generation cap of $P_{\text{gen}}=2400\,\text{kW}$ is applied to both feeders.
-Hyperparameters ($\gamma$, $\lambda$, learning rates, batch sizes, etc.) remain identical across systems unless scale adjustments are required.
-IV-C Hardware and Implementation Details
-Training is performed on a workstation with an NVIDIA RTX 6000 Ada GPU (48 GB), AMD Threadripper PRO CPU, and 256 GB RAM.
-OpenDSS runs on the CPU, while PyTorch handles actor–critic updates on the GPU.
-The HARL framework supports decentralized actors, a centralized critic, sequential PPO updates, and synchronized rollout collection.
-IV-D Evaluation Metrics
-Performance is evaluated using cumulative reward (weighted restoration benefit), restored active power (relative to the 2400 kW cap), convergence stability across seeds, inference latency per switching action, and total training time.
-This unified protocol enables consistent comparison across feeders and baseline RL/MARL methods.
+All experiments follow the unified HARL workflow in Section III (Algorithm 1).
+Each $T$-step episode uses decentralized recurrent actors for discrete switching actions and a centralized critic $V_{\phi}(s_{t})$ for values and $\lambda$-GAE advantages.
+Policy learning employs PPO-style updates with clipping threshold $\epsilon$, entropy coefficient $\beta_{\text{ent}}$, and sequential per-agent optimization.
+Rewards combine incremental restored power, normalized loss penalties, and continuous constraint violations.
+Both feeders use a fixed generation cap $P_{\text{gen}}=2400\,\text{kW}$.
+Hyperparameters ($\gamma$, $\lambda$, learning rates, batch sizes) remain identical unless scale adjustments are needed.
+IV-C Hardware and Evaluation Metrics
+Training is performed on a workstation with an NVIDIA RTX 6000 Ada GPU (48 GB), AMD Threadripper PRO CPU, and 256 GB RAM, with OpenDSS on CPU and PyTorch on GPU.
+Performance is evaluated using restored active power, convergence stability across seeds, inference latency, and training time, enabling consistent comparison across feeders and baseline RL/MARL methods.
 V Numerical Results
-This section evaluates the proposed HAPPO-based restoration framework on the IEEE 123-bus and IEEE 8500-node systems using the unified HARL training pipeline described in Sections IIIIV.
-Performance is reported in terms of (i) cumulative reward, which reflects the weighted restoration benefit accumulated during each episode, and (ii) restored active power, which directly measures physical restoration effectiveness relative to the system-wide generation cap $P_{\text{gen}}=2400\,\text{kW}$.
-All results are averaged over multiple independent random seeds.
-V-A Training Performance on IEEE 123-Bus and IEEE 8500-Node Systems
-Figures 3 and 4 show the evolution of cumulative reward and restored active power during training for the two feeders.
+V-A Performance on IEEE 123- and 8500-Bus Systems
+Figures 3 and 4 show the evolution of restored active power during training for the two feeders.
 Each curve corresponds to an independent seed initialized with randomized outage locations, critical-load placements, and DER conditions.
-On both feeders, agents select among three discrete switching actions (open, close, or no-op) and receive a shared reward that incorporates incremental restored power, normalized loss penalties, and differentiable constraint-violation terms.
-Unlike masking-based MARL methods, the proposed penalty-driven reward ensures uninterrupted learning near feasibility boundaries and supports smooth gradient propagation.
-In the IEEE 123-bus feeder 5 agents coordinate 26 switches.
-Across five independent training runs, the cumulative reward increases steadily and converges to $(3.6\pm 2.9)\times 10^{6},$ while the restored power stabilizes at $2294\pm 52~\text{kW}\quad(95.6\%\pm 2.2\%\text{ of }P_{\text{gen}}).$ As shown in Fig.
+In the IEEE 123-bus feeder, 5 agents coordinate 26 switches.
+Across five independent random seeds, the restored power stabilizes at $2294\pm 53~\text{kW}(95.6\%\pm 2.2\%\text{ of }P_{\text{gen}}).$ As shown in Fig.
 3, performance across seeds is tightly clustered, indicating reliable convergence and stable sequential actor updates.
@@
 Despite the dramatically larger state–action space and deeper network topology, HAPPO maintains monotonic improvement.
-The cumulative reward converges to $(2.9\pm 3.3)\times 10^{6},$ and restored power reaches $2308\pm 55~\text{kW}\quad(\!96.2\%\pm 2.3\%\text{ of }P_{\text{gen}}),$ as shown in Fig.
+The restored power reaches $2309\pm 55~\text{kW}\quad(\!96.2\%\pm 2.3\%\text{ of }P_{\text{gen}}),$ as shown in Fig.
 4.
 These results highlight the scalability of sequential trust-region updates, centralized advantage estimation, and penalty-based reward shaping.
-Figure 3: Training performance of HAPPO on the IEEE 123-bus feeder across four random seeds. Left: cumulative reward. Right: restored active power.
-Figure 4: Training performance of HAPPO on the IEEE 8500-node feeder across five random seeds. Left: cumulative reward. Right: restored active power.
-TABLE I: Performance comparison of HAPPO with state-of-the-art RL/MARL methods on IEEE 123-bus and IEEE 8500-node feeders (generation cap $P_{\text{gen}}=2400\,\text{kW}$)
+Figure 3: Training performance of HAPPO on the IEEE 123-bus feeder across five random seeds showing restored active power evolution.
+Figure 4: Training performance of HAPPO on the IEEE 8500-bus system across three random seeds showing restored active power evolution.
+TABLE I: Performance comparison of HAPPO with state-of-the-art RL/MARL methods on IEEE 123-bus and IEEE 8500-node feeders
 Method IEEE 123 IEEE 8500
-Cum. Reward Restored (%) Train Time (min) Inference (ms) Cum. Reward Restored (%) Train Time (min) Inference (ms)
-DQN [13] 28.38 $\pm$ 0.14 73.90 $\pm$ 0.46 34.7 4.2 65.58 $\pm$ 0.25 69.50 $\pm$ 0.28 249.3 6.8
-PPO [27] 29.01 $\pm$ 0.15 74.85 $\pm$ 0.47 44.8 7.1 67.94 $\pm$ 0.35 71.05 $\pm$ 0.27 255.7 9.8
-MAES [17] 29.78 $\pm$ 0.19 76.14 $\pm$ 0.40 91.8 25.4 69.82 $\pm$ 0.32 72.81 $\pm$ 0.23 307.0 38.7
-MAGDPG [18] 30.08 $\pm$ 0.17 77.00 $\pm$ 0.42 118.2 29.5 70.56 $\pm$ 0.24 73.44 $\pm$ 0.32 321.5 45.6
-MADQN [24] 30.58 $\pm$ 0.22 77.80 $\pm$ 0.59 98.8 12.6 71.37 $\pm$ 0.34 74.02 $\pm$ 0.39 279.2 19.0
-Mean-Field RL [19] 31.56 $\pm$ 0.25 80.26 $\pm$ 0.52 108.3 27.3 73.46 $\pm$ 0.33 77.40 $\pm$ 0.38 303.9 42.6
-QMIX [20] 33.38 $\pm$ 0.25 83.25 $\pm$ 0.45 96.4 15.8 76.62 $\pm$ 0.34 80.10 $\pm$ 0.30 291.3 23.5
-[…]
+Restored (%) Train Time (min) Inference (ms) Restored (%) Train Time (min) Inference (ms)
+DQN [13] 73.90 $\pm$ 0.46 34.7 4.2 69.50 $\pm$ 0.28 249.3 6.8
+PPO [27] 74.85 $\pm$ 0.47 44.8 7.1 71.05 $\pm$ 0.27 255.7 9.8
+MAES [17] 76.14 $\pm$ 0.40 91.8 25.4 72.81 $\pm$ 0.23 307.0 38.7
+MAGDPG [18] 77.00 $\pm$ 0.42 118.2 29.5 73.44 $\pm$ 0.32 321.5 45.6
+MADQN [24] 77.80 $\pm$ 0.59 98.8 12.6 74.02 $\pm$ 0.39 279.2 19.0
+Mean-Field RL [19] 80.26 $\pm$ 0.52 108.3 27.3 77.40 $\pm$ 0.38 303.9 42.6
+QMIX [20] 83.25 $\pm$ 0.45 96.4 15.8 80.10 $\pm$ 0.30 291.3 23.5
+HAPPO (ours) 95.6 $\pm$ 2.2 110.5 22.1 96.2 $\pm$ 2.3 391.2 33.4
 V-B Multi-Seed Convergence and Reproducibility
@@
 3 and 4 exhibit consistent monotonic convergence in both metrics.
-On the IEEE 123-bus feeder, restored power converges within $2230\text{--}2367~\text{kW},$ corresponding to 93–99% of the generation cap.
-On the IEEE 8500-node feeder, restored power stabilizes near $2224\text{--}2347~\text{kW},$ representing 93–98% of the available capacity.
+On the IEEE 123-bus feeder, restored power converges within $2241\text{--}2347~\text{kW},$ corresponding to 93–98% of the generation cap.
+On the IEEE 8500-node feeder, restored power stabilizes near $2254\text{--}2364~\text{kW},$ representing 94–99% of the available capacity.
 These results confirm that the proposed HARL-based HAPPO framework provides highly reproducible performance due to (i) sequential per-agent trust-region updates, (ii) low-variance centralized $\lambda$-GAE advantage estimation, and (iii) penalty-driven constraint handling that avoids disruptive early terminations.
@@
 Table I summarizes quantitative performance.
-HAPPO achieves the highest restored power and cumulative reward on both feeders.
+HAPPO achieves the highest restored power on both feeders.
 Its sequential per-agent trust-region updates and centralized advantage estimation enable stable policy optimization even in the presence of strong physical coupling and high-dimensional state spaces, outperforming value-based and value-decomposition methods that struggle under these conditions.
@@
 These findings confirm that the HARL-based HAPPO framework is well suited for large-scale, constraint-driven distribution system restoration.
+Although a full ablation study is beyond this work’s scope, benchmark methods lacking centralized critics or heterogeneous-agent modeling (e.g., independent PPO, parameter-sharing MAPPO) show reduced stability in strongly coupled feeders.
+The consistent multi-seed convergence in Figures 34 suggest centralized advantage estimation and microgrid-level heterogeneity are key contributors; systematic ablation remains for future work.
+Training uses full AC power-flow simulation on a single GPU, as offline training is not time-critical.
+At deployment, policy inference runs in milliseconds (under 35 ms per step on the IEEE 8500-node feeder), meeting practical restoration requirements.
 VI Conclusion
-This paper presented a HAPPO approach for distribution system restoration within a HARL framework using an OpenDSS-based multi-agent environment.
-Sequential trust-region updates combined with centralized advantage estimation enable stable, sample-efficient coordination among structurally heterogeneous microgrid agents, avoiding the instability issues common in conventional MARL.
-A continuous penalty-based reward replaces masking or early termination, yielding smooth gradients and naturally constraint-aware switching behavior.
-Experiments on the IEEE 123-bus and IEEE 8500-node feeders show that HAPPO consistently outperforms PPO, QMIX, Mean-Field RL, and other baselines.
-Under the 2400 kW generation cap, it restores roughly 92% and 96% of available load on the two systems, respectively, with strong multi-seed reproducibility and low inference latency.
-The OpenDSS integration ensures physical validity and supports direct transition to operational settings.
-Overall, the results demonstrate that heterogeneous-agent policy optimization provides a scalable, stable, and realistic solution for resilient distribution system restoration.
-Future work will incorporate model-based elements, communication constraints, partial observability, and dynamic DER participation to further strengthen real-time applicability.
+This paper presented a HAPPO-based HARL framework for distribution system restoration using OpenDSS.
+Sequential trust-region updates with centralized advantage estimation enable stable coordination among heterogeneous microgrid agents, while continuous penalty-based rewards provide constraint-aware switching behavior.
+Experiments on the IEEE 123-bus and 8500-node feeders show that HAPPO outperforms DQN, PPO, MAES, MAGDPG, MADQN, Mean-Field RL, and QMIX, restoring over 95% of available load under a 2400 kW cap with strong multi-seed stability and under 35 ms inference latency.
+Future work will explore model-based extensions, communication limits, partial observability, and dynamic DER participation.v
 Acknowledgment